From mar@roe.ac.uk Fri Aug 24 15:41:55 2012 Date: Fri, 24 Aug 2012 15:31:49 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Andrew Lawrence , Bob Mann , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Norman Gray , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives project team meeting minutes, 24/8/2012 Present: MAR, RSC, ETWS, MSH, RPB, RGM Apologies: JPE, AL, STV, NG, NJC, KTN, NCH DoNM: 10am Friday Sept. 7th 2012 in the Vista Hut (subject to non-demolition) NCH in the chair. Actions discharged this time: ----------------------------- None Actions partly discharged but continuing: ----------------------------------------- ACTION: RSC to scope out the areas/situations where we'd like to test MonetDB i.e. where do the current bottlenecks lie? Continues; the aim is to progress this early next week so that RGM can follow-up with the MonetDB folk. ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; (unless NJC has had the chance) ACTION: NCH to check the GCS for pawsky-related issues, and liaise with SJW over the LAS results NCH has started looking at this but more investigations are required. Actions carried forward from 10/08/12 meeting --------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. Specific points and new actions: -------------------------------- Project management: RGM reported that, as expected, the grant renewal would be partially held up by the Programmatic Review. Details are yet to emerge. WFCAM, VISTA and VST updates: Nothing to report this time. Comments and issues arising from CASU minutes: No new minutes this time Networking: Transfers are up-to-date WSA/VSA/OSA/GES Operations: ETWS noted: -- Catalogue data ingest for the OSA has completed, but some days need to be checked for broken files. -- VVV catalogue ingests for P88 are nearly done. -- Working on repopulating the missing FITS keywords due to the pyfits bug. Hardware and Systems: Ramses12 is broken, Eclipse were due to come in later in the day. The potential hosting of SDSS DR9 would require some creative disk juggling or possibly some new hardware. Survey Data Release: UKIDSS DR10 still awaiting QC and pawsky investigations. Feedback on the phase 3 catalogue submissions had been received from ESO via the VMC and VIDEO PIs. There were several issues that ESO want addressed. NJC has modded the code accordingly though we are still seeking clarification on a few of the points raised. The main sticking point currently is the requirement for IAU names to have a maximum of 2 decimal places AND be unique. Although the VVV data, recently prepared for ESO, would suffer from most of the issues raised, the VVV PIs had requested it still be uploaded to ESO, and this had been done. The recent issues with the ordering of deep stacking/tiling, missing header keywords (see below) and ramses12 have stalled the VMC and VHS proprietary releases. Software: RSC reported: NJC has modified CU13 to do deep stacking of OB frames in the correct order to fix the photometry issue in deep VVV tiles. Discovered yet another PyFITS 3.x bug that has lead to 10,000 OB files from VVV, VMC & WFCAM non-surveys losing important header keywords upon recalibration (following deep stacking) since the beginning of the year (when we upgraded to the PyFITS 3.x branch). This bug is fixed in the PyFITS 3.0.9 version that was released earlier this month, and Eckhard is running a script to fix the broken files using the values that were safely stored in the database prior to the corruption of the files. Additional modifications to our EsoRelease script based on the latest feedback from ESO including: a new FITS table column header key to denote indexed columns, a new schema tag to provide shorter descriptions of table columns, fettling of the EPS_REG keyword values for VIDEO and a fix for missing MAGLIM values. MAR noted a small mod to the UI to allow WSA non-survey users access to calibration files. ACTION: MAR to add notes on the VSA known issues page relating to the problems with deep stack photometry in VVV. Non-survey Data Release: None Astrogrid deployment & Data Analysis services: MSH reported the migration of Astrogrid services to ROE continues. Miscellaneous: NJC's talk at the ESO workshop in October had been accepted. MAR would attend the VHS meeting at the IoA in September. RGM was off to join NJC at the IAU meeting in Beijing. RSC and MSH had holidays scheduled over the next few weeks. -- Scanned by iCritical.